A 1.26⋅10−21.26⋅10-2 M solution of NaOH is made.

What is the pOH of this solution?

What is the pH of this solution?

What is the [H+] in the solution? Record this answer in scientific notation using 3 sig figs.
Answer #1

Answer

The ionization reaction of NaOH is

NaOH(aq) ------> Na+(aq)   + OH- (aq)

it is a strong base it completely dissociates into ions

so concentration of OH- is equal to concentration of NaOH

[OH-] = [NaOH]= 1.26*10^-2 M

We know that

pOH= -log[OH-] = -log[1.26*10^-2]= -[-1.89] =1.89

We know that

pH+pOH= 14 or pH= 14-pOH

hence pH= 14-1.89 = 12.1

and

[H+]= 10^-pH = 10^-12.1 = 7.94*10^-13 M